# SASG-DA: Sparse-Aware Semantic-Guided Diffusion Augmentation For Myoelectric Gesture Recognition

### v2

@@ 6.1 The Faithfulness of Generated Samples
Figure 5: Distribution visualizations of the original (blue) and generated (red) samples for subject 1 from (a) DB7, (b) DB4 and (c) DB2. Different shades of each color denote different gesture classes.
Figure 6: Examples of original (blue) and generated (red) samples of (a) DB7, (b) DB4 and (c) DB2.
To verify the distributional faithfulness of the generated samples, we randomly select a subject from DB7, DB4, and DB2 for analysis. For each subject, features are extracted from both real and generated samples using a pretrained classifier, and subsequently projected into a two-dimensional space via t-SNE. As shown in Fig. 5, the generated samples consistently align with the corresponding real samples in the embedding space, demonstrating a high degree of distributional similarity. These results indicate that the proposed Semantic Representation Guidance mechanism effectively preserves the intrinsic structure of the original data, facilitating the generation of class-consistent samples within the real data manifold.
We further provide qualitative comparisons by randomly selecting one sample from each dataset. As illustrated in Fig. 6, the generated sEMG signals exhibit waveform patterns and temporal dynamics that closely resemble those of the corresponding real samples. In particular, even fine-grained signal details, such as subtle fluctuations and transient peaks, are well generated, indicating that the model is capable of generating high-fidelity sEMG signals with refined structure.
6.2 Extension of Dataset Coverage
To verify that our SASS strategy explicitly expands the data distribution, we conduct condition space visualization and quantitative sparsity evaluations. Specifically, we use t-SNE to project the semantic representations of the original training set, as well as those obtained through the GMSS and SASS mechanisms, into a two-dimensional space. As shown in Fig. 7, the semantic representations generated by the GMSS mechanism are evenly distributed within the original feature space, producing diverse and distribution-consistent conditions. Furthermore, the semantic representations generated by the SASS mechanism are distributed in the sparser regions of the original feature space, which significantly enlarges the coverage of the training distribution while maintaining consistency with the original data manifold.
We further verify that the samples generated under sparse condition guidance are indeed sparse in nature through quantitative metrics. Specifically, following the settings in [sehwag2022generating, um2024self], we adopt three sparsity measures on generated samples: the average k-nearest neighbor distance (AvgKNN), the Local Outlier Factor (LOF), and the Rarity Score. As shown in Fig. 8, the generated samples guided by the sparse conditions consistently exhibit higher AvgKNN distances, larger LOF values, and higher Rarity Scores compared to those generated by the GMSS. These results quantitatively demonstrate that the SASS mechanism successfully shifts the generated samples towards the sparse regions of the data distribution, explicitly expanding the coverage of the training manifold. This property is particularly beneficial for enhancing the generalization capability of the classifier by exposing it to more diverse and underrepresented regions of the original data space.
@@ 6.2 Extension of Dataset Coverage
Table 8: Performance comparison of different guidance strategies on DB7 and DB4.
Method DB7 DB4
FID CAS Crossformer TDCT STCNet FID CAS Crossformer TDCT STCNet
Label Condition 2.77 49.83 80.49 77.09 80.62 2.95 45.99 68.81 66.04 68.19
Classifier Guidance 1.73 86.08 79.72 76.82 80.11 2.03 92.18 67.81 65.23 68.21
Classifier-Free Guidance 1.62 77.00 80.54 77.46 80.95 2.13 71.44 68.91 66.30 68.43
GMSS 0.88 71.73 80.87 77.57 81.67 0.87 68.09 69.44 66.99 69.41
SASG-DA 1.35 58.65 81.31 78.77 82.15 1.37 56.99 69.73 67.74 70.05
@@ 6.3 Trade-off Between Faithfulness and Diversity
To illustrate the trade-off between faithfulness and diversity, we compare our method with three common conditional generation baselines: Label Condition, Classifier Guidance [dhariwal2021diffusion], and Classifier-Free Guidance [ho2021classifier]. The results are shown in Table 8. The Label Condition approach introduces only coarse-grained class information, which results in relatively high FID scores (e.g., 2.77 on DB7) and low CAS scores (49.83), indicating insufficient faithfulness to the target classes. In contrast, Classifier Guidance employs an explicit classifier to steer the generation toward class-consistent samples [dhariwal2021diffusion], significantly improving CAS (e.g., up to 86.08) and reducing FID. However, stronger classifier gradient guidance can overly constrain the generation process, limiting the diversity of the synthesized samples by aligning them too closely with the classifier’s decision boundaries. Classifier-Free Guidance avoids the need for an external classifier by enabling the generative model to produce both conditional and unconditional outputs during inference [ho2021classifier]. While it avoids the overhead of an external classifier, stronger condition strength still tends to reduce diversity by pulling generated samples toward the high-density regions of the class distribution.
@@
6.4 Effectiveness in Mitigating Overfitting
To evaluate the effectiveness of our method in mitigating overfitting, we analyze the training and testing accuracy curves on randomly selected samples from DB4 and DB7. As shown in Fig. 9, the baseline model exhibits a noticeable divergence between training and testing accuracy as training progresses, indicating a tendency to overfit the training data. In contrast, our method effectively narrows the gap, demonstrating enhanced generalization and mitigated overfitting. This indicates that our data augmentation strategy enhances model generalization by providing the model with more diverse and effective samples, thereby reducing its tendency to overfit to spurious patterns or redundant data.
Figure 9: Training and testing accuracy curves for randomly selected subjects from (a) DB7 and (b) DB4.
6.5 Size and Diversity of Synthetic Data
@@
We further explore the effect of the size and diversity of the synthetic samples on model performance. To this end, we systematically increase the size of the synthetic dataset to 1×, 2×, 3×, and 4× the size of the original training set. As shown in Fig. 10, experiments on both DB7 and DB4 demonstrate that the classification accuracy improves monotonically with the growth of synthetic data. Specifically, for DB7, we observe consistent gains from +8.54% to +9.88% compared to the baseline when enlarging the synthetic set. A similar trend is seen on DB4, where the accuracy steadily rises from +8.80% to +9.92% above the baseline. These results suggest that expanding the size and diversity of synthetic samples contributes to more robust model generalization.
@@ 6.6 Limitation and Future Work
Although the proposed diffusion-based augmentation method effectively improves classification performance by generating faithful and diverse samples, its generation efficiency remains a limitation [yin2024one]. Due to the iterative nature of the diffusion process, producing a large amount of synthetic data can be computationally expensive and time-consuming compared to other generative models. In future work, we plan to investigate acceleration strategies, such as distillation-based fast samplers [yin2024one], to reduce generation costs without compromising sample diversity and fidelity. Exploring lightweight diffusion architectures could further facilitate practical deployment in data augmentation applications.
@@ 7 Conclusion
Informative data scarcity and overfitting remain major challenges for deploying sEMG-based gesture recognition in real-world human–machine interaction (HMI) applications. To address these issues, we propose a novel diffusion-based data augmentation framework, SASG-DA, which integrates semantic representation guidance and sparse-aware sampling to generate both faithful and diverse training samples. By modeling and selectively sampling from a learned semantic embedding space, our method enables flexible and controllable augmentation that not only enhances class fidelity but also improves coverage of underrepresented regions. Extensive experiments on multiple public sEMG datasets demonstrate that SASG-DA consistently improves recognition accuracy and generalization over existing augmentation strategies. We believe this work provides a practical and principled data augmentation approach for improving gesture recognition performance. Moreover, the proposed framework holds promise for broader application in biosignal and time-series classification tasks where data scarcity and overfitting are common challenges. Future work will explore extending this augmentation approach to other domains where semantically conditioned and diversity-aware sample generation is beneficial.

### v3

@@ 6.1 The Faithfulness of Generated Samples
To assess distributional faithfulness, we randomly select one subject from DB7 and DB4. Features are extracted from both real and generated samples using a pretrained classifier, and subsequently projected into a two-dimensional space via t-SNE. As shown in Fig. 5, generated samples closely align with real ones in the embedding space, indicating high distributional similarity. This demonstrates that the proposed Semantic Representation Guidance mechanism effectively preserves the intrinsic data structure and generates class-consistent samples within the real data manifold.
Figure 5: Distribution visualizations of the original (blue) and generated (red) samples for subject 1 from (a) DB7 and (b) DB4. Different shades of each color denote different gesture classes.
Figure 6: Examples of original (blue) and generated (red) samples of (a) DB7 and (b) DB4.
We further provide qualitative comparisons by randomly selecting one sample from each dataset. As illustrated in Fig. 6, the generated sEMG signals exhibit waveform patterns and temporal dynamics that closely resemble those of the corresponding real samples. In particular, even fine-grained signal details, such as subtle fluctuations and transient peaks, are well generated, indicating that the model is capable of generating high-fidelity sEMG signals with refined structure. Further analyses of the physiological plausibility of the synthetic data are presented in the Supplementary Material.
@@
6.2 Extension of Dataset Coverage and Downstream Utility
To verify that our SASS strategy explicitly expands the data distribution, we conduct condition space visualization, quantitative sparsity evaluations and downstream utility analysis. Specifically, we use t-SNE to project the semantic representations of the original training set, as well as those obtained through the GMSS and SASS mechanisms, into a two-dimensional space. As shown in Fig. 7, the semantic representations generated by the GMSS mechanism are evenly distributed within the original feature space, producing diverse and distribution-consistent conditions. Furthermore, the semantic representations generated by the SASS mechanism are distributed in the sparser regions of the original feature space, which significantly enlarges the coverage of the semantic space while maintaining consistency with the original data manifold.
We further verify that the samples generated under sparse condition guidance are indeed sparse in nature through quantitative metrics. Specifically, following the settings in [24, 25], we adopt three sparsity measures on generated samples: the average k-nearest neighbor distance (AvgKNN), the Local Outlier Factor (LOF), and the rarity score. As shown in Fig. 8, the generated samples guided by the sparse conditions consistently exhibit higher AvgKNN distances, larger LOF values, and higher rarity scores compared to those generated by the GMSS. These results quantitatively demonstrate that the SASS mechanism successfully shifts the generated samples towards the sparse regions of the data distribution, explicitly expanding the coverage of the training manifold. This property is particularly beneficial for enhancing the generalization capability of the classifier by exposing it to more diverse and underrepresented regions of the original data space.
Figure 9: Downstream utility validation by (a) Performance trend across sparsity levels and confidence for generated samples. (b) Training dynamics for a randomly selected subject from DB7.
Beyond overall classification performance, we further validate the downstream utility in generating targeted diverse samples by performing sparsity-stratified training analysis. Specifically, we mix samples generated by GMSS and SASS to form a broad distributional range, and then divide them into five equally sized levels based on their rarity scores. Level 0 represents the baseline without augmentation. We evaluate each sparsity level on DB7 and DB4. The results in Fig. 9 show that downstream performance consistently improves as sparsity increases and stabilizes at a high level, indicating that the SASS mechanism effectively enhances the utility of generated samples for downstream tasks by enhancing their targeted diversity. While the performance gain gradually saturates at Level 5, highly sparse samples remain valid and do not cause noticeable performance degradation. This observation is consistent with our motivation of explicitly expanding the training distribution to improve downstream utility.
We additionally analyze generated sample confidence across sparsity levels, showing that confidence decreases as sparsity increases, consistent with samples from underrepresented regions being less confidently predicted by the pretrained classifier on the original training set. Furthermore, the training curves in Fig. 9 show that our approach narrows the gap between training and test accuracy compared to the baseline, suggesting mitigated overfitting and improved generalization.
@@ 6.3 Trade-off Between Faithfulness and Diversity
To illustrate the trade-off between faithfulness and diversity, we compare our method with three common conditional generation baselines: Label Condition, Classifier Guidance [5], and Classifier-Free Guidance [16]. The results are shown in Table 8. The Label Condition approach introduces only coarse-grained class information, which results in relatively high FID scores (e.g., 2.77 on DB7) and low CAS scores (49.83), indicating insufficient faithfulness to the target classes. In contrast, Classifier Guidance employs an explicit classifier to steer the generation toward class-consistent samples [5], significantly improving CAS (e.g., up to 86.08) and reducing FID. However, stronger classifier gradient guidance can overly constrain the generation process, limiting the diversity of the synthesized samples by aligning them too closely with the classifier’s decision boundaries. Classifier-Free Guidance avoids the need for an external classifier by enabling the generative model to produce both conditional and unconditional outputs during inference [16]. While it avoids the overhead of an external classifier, stronger condition strength still tends to reduce diversity by pulling generated samples toward the high-density regions of the class distribution.
Table 8: Performance comparison of different guidance strategies on DB7 and DB4. FID and cFID values are reported with cFID computed conditionally per class. For clarity, cFID A-T refers to the Augmented training set vs the Test set.
Method DB7 (cFID Train-Test=5.97) DB4 (cFID Train-Test=8.64)
FID CAS cFID A-T Crossformer TDCT STCNet FID CAS cFID A-T Crossformer TDCT STCNet
Label Condition 2.77 49.83 5.15 80.49 77.09 80.62 2.95 45.99 8.24 68.81 66.04 68.19
Classifier Guidance 1.73 86.08 4.83 79.72 76.82 80.11 2.03 92.18 7.38 67.81 65.23 68.21
Classifier-Free Guidance 1.62 77.00 4.97 80.54 77.46 80.95 2.13 71.44 7.15 68.91 66.30 68.43
GMSS 0.88 71.73 4.27 80.87 77.57 81.67 0.87 68.09 5.75 69.44 66.99 69.41
SASG-DA 1.35 58.65 4.23 81.31 78.77 82.15 1.42 56.28 5.73 69.73 67.74 70.05
Table 9: Comparison of computational cost, method characteristics, and performance gain across datasets.
Method FLOPs (G) Params (M) Training Time (s/it) Inference Throughput (samples/s) Faithfulness Diversity Perf Gain (DB7 / DB4 / DB2)
E-TRGAN 0.51 2.28 0.02 476.3 $\checkmark$ 0.54 / 0.68 / 0.48
DiffMix 2.17 3.60 0.23 13.1 $\checkmark$ 0.57 / 0.60 / 0.10
SSE-LDM 0.40 1.36 0.08 21.0 $\checkmark$ 0.45 / 0.85 / 0.19
EEG-Enhancing 2.17 3.60 0.20 12.9 $\checkmark$ 0.24 / 0.47 / 0.26
NTD 0.37 0.56 0.04 25.1 $\checkmark$ 2.42 / 2.74 / 1.43
PatchEMG 4.34 3.60 0.38 7.6 $\checkmark$ 3.23 / 3.33 / 2.20
DESAM 2.11 2.18 0.21 10.48 $\checkmark$ 3.04 / 1.98 / 1.67
Minority 2.17 3.60 0.23 9.8 $\checkmark$ 3.24 / 3.32 / 2.45
CADS 2.17 3.60 0.23 12.8 $\checkmark$ 4.07 / 4.10 / 3.12
SASG-DA (500) 2.26 3.60 0.25 12.7 $\checkmark$ $\checkmark$ 5.70 / 5.96 / 4.73
SASG-DA (200) 2.26 3.60 0.25 32.4 $\checkmark$ $\checkmark$ 5.30 / 5.85 / 4.38
SASG-DA (50) 2.26 3.60 0.25 129.3 $\checkmark$ $\checkmark$ 5.17 / 5.67 / 4.36
@@
To further examine whether sparsely generated samples plausibly extend the training distribution, we analyze distributional coverage with respect to a held-out test set. Specifically, we compute cFID (class-wise FID), which averages per-class distances to better capture class-level alignment between augmented and test samples. We evaluate cFID between (i) the original training and test sets, and (ii) the augmented training set (Train + Synthetic) and the test set. As shown in Table 8, incorporating sparse samples consistently reduces the cFID gap, indicating that the generated samples expand the data manifold toward a more plausible distribution.
6.4 Overall Evaluation on Computational Efficiency and Performance
To provide a comprehensive evaluation of the proposed method, we compare the computational cost, method characteristics, and performance gains of generative augmentation approaches in Table 9. For fairness, all methods are evaluated under the same settings described in Section 5.3. FLOPs are reported per diffusion step or forward, training time per iteration depends on the batch size, and inference throughput varies with the number of sampling steps. Here, we focus on the relative computational overhead introduced by our method.
Compared with SOTAs built upon standard diffusion models (e.g., DiffMix [7], CADS [23]), our approach introduces only limited additional computational overhead while yielding consistent and notable performance improvements across three datasets. Importantly, it explicitly enhances faithfulness and diversity at minor cost, without relying on the costly CFG [16] or other gradient-based guidance [25]. Methods such as SSE-LDM [18] and NTD [20] achieve low computational cost through specially designed lightweight architectures; however, their performance improvements remain relatively modest. Compared with GAN-based approaches, our method benefits from the inherent stability and strong generative capacity of diffusion models, resulting in more substantial gains in downstream tasks. In addition, we evaluate the efficiency–performance trade-off under reduced sampling steps. Notably, fewer sampling steps substantially increase inference throughput while causing only minor performance loss, showing that our method remains effective in more efficient settings and holds the potential for practical applications. In practice, SASG-DA performs offline augmentation with a one-time generation cost of approximately 20, 7, or 2 minutes per subject for 500, 200, or 50 steps, respectively, generating around 15k samples on large-scale Ninapro datasets, without requiring additional generative computation during downstream training. Compared with single-sample augmentation strategies, SASG-DA provides more consistent gains across backbones and datasets, particularly under cross-subject settings, highlighting a practical trade-off where moderate offline cost yields improved generalization through continuous and reusable high-quality synthetic samples.
In addition, SASG-DA consistently improves performance across all three datasets and multiple backbone architectures, with average gains of 5.70%, 5.96%, and 4.73%, respectively. The variations in improvement are expected given differences in signal characteristics, gesture complexity, and baseline performance, yet the results clearly demonstrate the effectiveness of SASG-DA across diverse data characteristics.
6.5 Cross-subject Evaluation
To further evaluate the adaptability of the proposed SASG-DA, we conduct a cross-subject evaluation on the GrabMyo dataset [42]. As a relatively recent wrist-based dataset, it offers a representative and practical setting for cross-subject evaluation, consistent with current trends in wearable sEMG [1] acquisition and real-world usage scenarios. It includes sEMG signals from 43 subjects performing 17 gestures across 3 sessions, each comprising 7 trials recorded under similar conditions using 6 channels at 2048 Hz. We select the first session for evaluation and adopt a six-fold cross-subject cross-validation strategy, where subjects were sequentially divided into six groups. Before training, the raw sEMG signals were preprocessed using the official GrabMyo processing pipeline, where the data were segmented into 200 ms windows with a 150 ms step size. Since Crossformer [33] did not produce stable results on this dataset, we adopted GengNet [43] as a reliable backbone for downstream evaluation.
The results in Table 10 show that the proposed SASG-DA consistently improves recognition performance under the challenging cross-subject setting, outperforming all SOTA methods across three backbones (effect size $r>0.5$). While several augmentation techniques offer limited or even negligible gains in this scenario, SASG-DA provides clear and stable improvements, indicating that generating faithful and diverse samples effectively facilitates generalization to unseen subjects. These results highlight the robustness and adaptability of our method and its potential in practical applications.
Table 10: Performance comparison in terms of ACC (%) on dataset GrabMyo under cross-subject evaluations. The results report the mean and the standard deviation across folds.
Aug. Method TDCT STCNet GengNet
[…]
6.6 Sensitivity Analysis to Synthetic Data Size
@@
We further explore the effect of the augmented size and sensitivity of the synthetic samples on model performance. To this end, we systematically increase the size of the synthetic dataset to 0.5×, 1×, 2×, 3×, and 4× the size of the original training set. As shown in Fig. 10, experiments on both DB7 and DB4 demonstrate that the classification accuracy improves monotonically with the growth of synthetic data. Specifically, for DB7, we observe consistent gains from +6.45% to +9.88% compared to the baseline when enlarging the synthetic set. A similar trend is seen on DB4, where the accuracy steadily rises from +6.95% to +9.92% above the baseline. These results suggest that expanding the size and diversity of synthetic samples contributes to more robust model generalization. Furthermore, the consistent improvement indicates low sensitivity to the augmentation ratio, suggesting that our method preserves faithfulness to the real data distribution while providing beneficial diversity, without noticeable performance degradation within the tested range.
@@ 6.7 Limitation and Future Work
Although the proposed diffusion-based augmentation method effectively improves classification performance by generating faithful and diverse samples, its generation efficiency remains a limitation [44]. Due to the iterative nature of the diffusion process, producing a large amount of synthetic data can be computationally expensive and time-consuming compared to other generative models. Accordingly, SASG-DA is formulated as an offline augmentation strategy and may be less suitable for scenarios requiring on-the-fly generation of large-scale data. In future work, we plan to investigate acceleration strategies, such as distillation-based fast samplers [44], to reduce generation costs without compromising sample diversity and faithfulness. Exploring lightweight diffusion architectures could further facilitate practical deployment in data augmentation applications.
@@ 7 Conclusion
Informative data scarcity and overfitting remain key challenges for real-world sEMG-based gesture recognition in HMI applications. To address them, we propose SASG-DA, a diffusion-based augmentation framework that integrates semantic representation guidance with sparse-aware semantic sampling to generate faithful and targeted diverse samples. By modeling and selectively sampling from a learned semantic space, SASG-DA enables flexible and controllable augmentation that enhances class faithfulness and expands coverage of underrepresented regions. Extensive experiments on multiple public sEMG datasets show consistent improvements in recognition accuracy and generalization over existing augmentation methods. We believe SASG-DA offers a practical and principled data augmentation approach for gesture recognition and holds potential for broader biosignal and time-series classification tasks with data scarcity. Future work will explore extending this augmentation approach to cross-domain settings and other domains where semantically conditioned and diversity-aware sample generation is beneficial, as well as testing its applicability in real-time prosthetic control scenarios.